OpenStudy (anonymous):

The legs of a right triangle are lengths x and x√3. The cosine of the smallest angle of the triangle is _____.

OpenStudy (anonymous):



OpenStudy (anonymous):

Find the hypotenuse: h^2 = x^2 + 3x^2 = 4x^2 -> h = 2x. cos A = x*v3/2x = V3/2 = cos Pi/6 A = ....